OpenStudy (anonymous):

what it the laplace transform of y'' + 6y' + 9y = 0; initial values of y(0) = -1 and y'(0) = 6

OpenStudy (anonymous):

yay! come on goku!

OpenStudy (ash2326):

ha ha, just a min

OpenStudy (anonymous):

haha, no rush

OpenStudy (ash2326):

We know that \[L(y'')=s(sL(y)-y(0))-y'(0)\] and \[L(y')=sL(y)-y(0)\] Let's use them here

OpenStudy (anonymous):

I'm down to....

OpenStudy (anonymous):

Y(s) = -s/(s+3)^2

OpenStudy (ash2326):

\[ y'' + 6y' + 9y = 0\] Let's take laplace transform both sides \[L(y'')+6L(y')+9L(y)=0\] \[(s^2L(y)-sy(0)-y'(0))+6sL(y)-6y(0)+9L(y)=0\] \[L(y)(s^2+6s+9)+s-6+6=0\] \[L(y)=\frac{-s}{(s+3)^2}\] Yeah you're correct:D

OpenStudy (anonymous):

so how does all of the equal -e^(-3t) + 3te^(-3t)

OpenStudy (ash2326):

You have to take inverse Laplace transform to get that

OpenStudy (anonymous):

i know, but how? I need to get a better grasp on the concepts

OpenStudy (anonymous):

it's where i get stuck most of the time

OpenStudy (ash2326):

Let me try

OpenStudy (ash2326):

If we have \[L(e^{at})=\frac{1}{s-a}\] and \[L(te^{at})=\frac{-d}{ds} (\frac{1}{s-a})\] or \[L(te^{at})=\frac{1}{(s-a)^2}\]

OpenStudy (ash2326):

We have \[L(y)=\frac{-s}{(s-3)^2}\] it can be written as \[L(y)=\frac{-s+3-3}{(s-3)^2}=\frac{-3}{(s-3)^2}-\frac{(s-3)}{(s-3)^2}\] Do you get this?

OpenStudy (anonymous):

damnit... i always forget to add to the numerator...

OpenStudy (anonymous):

well, at least make it look alike. do you have time to walk me through a harder one?

OpenStudy (ash2326):

Oops, I gotta go now. Sorry Mention me there like this @IStutts , I'll try to come and help

OpenStudy (anonymous):

thank you